OpenStudy (anonymous):

Trying to solve the equation: z^2+8z=-15

OpenStudy (solomonzelman):

\(\normalsize\color{darkgreen}{ \rm z^2+8z=-15 }\) ` Way 1. ` Add 16 to both sides, and complete the square. ` Way 2. ` Add 15 to both sides and factor. HINT: \(\normalsize\color{darkgreen}{ \rm 8=5+3 }\) ` Way 3. ` Quadratic formula.

OpenStudy (solomonzelman):

I prefer way 2

OpenStudy (solomonzelman):

@hrassistant , are you there ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

So try one of the ways... either one is fine, just that I suggest to do it the second way.

OpenStudy (anonymous):

so I take the equation and add 15 to both sides leaving the one side with a zero

OpenStudy (solomonzelman):

yes, exactly :)

OpenStudy (anonymous):

so now I have z^2+8z+15=0

OpenStudy (anonymous):

can I combine terms

OpenStudy (solomonzelman):

You can not add anything in the equation, but you can factor it.

OpenStudy (anonymous):

z(z+8)+15

OpenStudy (solomonzelman):

No

OpenStudy (solomonzelman):

\(\Large\color{darkgreen}{ (x+ \text{_})( x+ \text{_}) }\) fill in the blanks.

OpenStudy (anonymous):

(z+4)(z+2)

OpenStudy (anonymous):

oops

OpenStudy (solomonzelman):

4 and 2 ? You are incorrect again. Lets get back to factoring later, try to use the quadratic formula.

OpenStudy (anonymous):

(z+5)(z+3)

OpenStudy (solomonzelman):

Yes, ` (x+5)(x+3) ` is correct.

OpenStudy (solomonzelman):

So this is what you/we are doing.... \(\Large\color{darkgreen}{ z^2+8z=-15 }\) \(\Large\color{darkgreen}{ z^2+8z\color{red}{+15}=-15 \color{red}{+15} }\) \(\Large\color{darkgreen}{ z^2+8z+15=0 }\) \(\Large\color{darkgreen}{(z+3)(z+5)=0 }\)

OpenStudy (anonymous):

thank you!!!

OpenStudy (solomonzelman):

You welcome:) IF you want to know I can show how to complete the square, but you don't have to worry about it if you don't want to.